# Multistability of graphene nanobubbles

### v1

@@ Abstract
Using the example of Ar, Kr, and Xe atoms, it is shown that graphene nanobubbles on flat substrates are multistable systems. A nanobubble can have many stable stationary states, each characterized by the number of layers, $l$, within the cluster of internal atoms. The layers are circular in shape, concentrically stacked on top of each other, forming an $l$-stepped pyramid with a flat top. The covering of this pyramid with a graphene sheet is achieved through its local stretching. The valence bonds of the sheet stretch only over the group of internal atoms; outside the coverage zone, the sheet remains undeformed and lies flush against the substrate. The maximum possible number of layers, $l_{m}$, increases monotonically with the number of atoms $N$ ($l_{m}=6$ for $N=4000$). The graphene sheet, interacting with the substrate, compresses the internal atom cluster against it, generating an internal pressure of $P\sim 1$ GPa. Numerical simulations of thermal vibrations reveal that among all $l$-layer configurations of a nanobubble, there is always one ”ground” state. Upon heating, this ground state smoothly transitions into a layerless liquid state. All other stationary states transform into this ground state once a certain temperature is reached (for $N=4000$, the ground state corresponds to state with $l=4$). The coexistence of several stable states with different numbers of layers at low temperatures leads to the absence of a universal shape for the nanobubbles. In this scenario, the height-to-radius ratio, $H/R$, can vary from 0 to 0.24, depending on the number of layers.
Keywords: Van der Walls bubble, graphene nanobubble, molecular modelling, phase transition
@@ I Introduction
In addition to experimental methods, analytical models and numerical simulations are used to study nanobubbles. Analytical models employ elastic plate theory, taking into account the contribution of adhesion energy between the layers of van der Waals heterostructures Yue2012 ; Wang2013a ; Dai2018 . More complex models also incorporate the equation of state of the substance inside the nanobubble Zhilyaev2019 ; Iakovlev2019 ; Aslymov2020 ; Qu2023 and nonlinear effects Lyublinskaya2020 . Using such models, it has been shown that nanobubbles with a circular shape exhibit universal scaling behavior: the ratio of the nanobubble’s height to its radius, $H/R$, remains constant. Numerical simulations allow not only for the verification of analytical results but also for gaining deeper insights into nanobubble properties. It has been demonstrated that spatial confinement and internal pressure within a nanobubble promote the formation of internal crystalline structures, which can be planar in shape Iakovlev2017 , and that the nanobubble’s shape can critically depend on the properties of the encapsulated substance Ghorbanfekr2017 . Numerical simulations have shown Jain2017 that a nanobubble containing an ideal gas always has a circular shape with a constant ratio $H/R=0.204$. Molecular dynamics simulations have modeled transitions of argon inside a nanobubble into a liquid-crystal state Korneva2023 . It has been demonstrated that the presence of nanobubbles can lead to a reduction in the interfacial thermal conductivity of layered van der Waals structures Qu2023a , as well as cause their significant thermal contraction Qu2023b .
In the paper, all possible stationary states of graphene nanobubbles will be numerically modeled. This paper is organized as follows. Section II describes the model used. Section III presents the possible stationary states of the nanobubbles. The stability of these stationary states against thermal fluctuations is analyzed in Section IV. The main results are summarized in the Conclusion (Section V).
@@ II Model
Consider a rectangular graphene sheet in which the chains of valent bonds along the $x$-axis have a ”zigzag” structure with $N_{x}$ nodes, and along the $y$-axis—an ”armchair” structure with $N_{y}$ nodes (the sheet lies parallel to the $z=0$ plane). Let us take a sheet with free edges—see Fig. 1. The sheet consists of $N_{c}=N_{x}N_{y}/2-2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its length is $L_{x}=(N_{x}-1)r_{c}\sqrt{3}/2$ and its width is $L_{y}=(3N_{y}/4-1)r_{c}$ ($r_{c}=1.418$ Åis the C–C bond length). In all calculations, we will use a graphene sheet with $N_{x}=416$, $N_{y}=480$ (sheet dimensions: $50.963\times 50.906$ nm2, number of carbon atoms $N_{c}=99838$). We will assume that hydrogen atoms are attached to the edge carbon atoms, which participate in the formation of only two C–C bonds, forming united CH atoms with them of mass $M_{e}=13m_{p}$ (the mass of the internal sheet atoms is $M_{c}=12m_{p}$, where $m_{p}=1.66\times 10^{-27}$ kg is the proton mass).
@@
where $\cos\varphi=({\bf u}_{n}-{\bf u}_{k},{\bf u}_{l}-{\bf u}_{k})/r_{nk}r_{kl}$, and $\varphi_{0}=2\pi/3$ is the equilibrium valent angle. Parameters $\alpha=1.7889$ nm-1 and $\epsilon_{\varphi}=1.3143$ eV can be found from the small amplitude oscillations spectrum of the graphene sheet Savin2008 .
@@
where $\varepsilon_{0}$ is the interaction energy (adhesion energy) and $h_{0}$ is the equilibrium distance to the surface of the half-space. For a silicon dioxide (SiO2) substrate, the energy is $\varepsilon_{0}=0.074$ eV and the distance is $h_{0}=5$ Å  Koenig2011 .
Table 1: Values of the parameters of the interaction potential with a flat substrate (5) for atoms C, Ar, Kr, Xe.
C Ar Kr Xe
$\varepsilon_{0}$ (eV)  0.074  0.1406  0.17242  0.19462
$h_{0}$ (Å) 5.0 5.0 5.116 5.397
@@ C Ar Kr Xe
where $N=N_{c}+N_{g}$ is the total number of atoms in the two-component molecular system, the vectors $\{{\bf u}_{n}\}_{n=N_{c}+1}^{N}$ specify the coordinates of the gas atoms, and $M_{g}$ is the atomic mass ($M_{g}=39.96$, 83.80, 131.29 $m_{p}$ for Ar, Kr, and Xe atoms, respectively). The first term in the sum (6) represents the kinetic energy of the gas subsystem, the second term—the interaction energy of the gas atoms with the flat substrate (the values of the interaction potential parameters (5) for the gas atoms are given in Table 1). The last term,
@@
defines the interaction energy between the gas atoms, where $r_{nk}=|{\bf u}_{n}-{\bf u}_{k}|$ is the distance between atoms $n$ and $k$. The interaction between atoms is described by the Lennard-Jones (6,12) potential:
@@
The values of the interaction potential parameters for various pairs of atoms are given in Table 2.
Table 2: Values of the interaction potential parameters (8) for pairs of atoms CAr, CKr, CXe, ArAr, KrKr, XeXe Iakovlev2017 ; Rappe1992 ; Vogt2014 .
CAr CKr CXe ArAr KrKr XeXe
$\varepsilon_{1}$ (eV)  0.00525  0.00643  0.00724  0.010  0.015  0.019
$\sigma$ (Å) 3.3965 3.4965 3.7465 3.40 3.60 4.10
@@
Figure 2: Stationary $l$-layer states of encapsulated $N_{g}=2909$ argon atoms with $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). For clarity of presentation, only one half of the graphene sheet is shown. The graphene sheet is depicted in gray. For argon atoms, their vertical displacement is indicated by color (blue corresponds to the minimum displacement with coordinate $z=h_{0}$, red corresponds to the maximum displacement with $z=h_{0}+10$ Å).
Figure 3: Distribution of vertical displacements $p(z)$ for the stationary states of encapsulated $N_{g}=3936$ krypton atoms with number of layers $l=1$, …, 5.
@@ III Stationary States of Graphene Nanobubbles
Let $\{{\bf u}_{n}^{0}=(x_{n}^{0},y_{n}^{0},z_{n}^{0})\}_{n=1}^{N}$ be a solution to problem (12). The corresponding stationary state is characterized by its energy $E(\{{\bf u}_{n}^{0}\}_{n=1}^{N})$, the maximum value of the vertical displacements of the graphene sheet over the encapsulated group of atoms (the height of the graphene bulge)
@@
We will consider that a sheet atom participates in forming the upper surface of the nanobubble if its vertical displacement $\Delta z_{n}=z_{n}^{0}-h_{0}>h_{1}$, where the displacement threshold is $h_{1}=1.5$ Å. The center of the nanobubble then has coordinates
@@
Here, the second vertical displacement threshold is $h_{2}=2.5$ Å, and the distance to the center is $r_{n}=[(x_{n}^{0}-x_{c})^{2}+(y_{n}^{0}-y_{c})^{2}]^{1/2}$.
@@
Figure 4: Stationary $l$-layer states of a nanobubble with $N_{g}=3936$ krypton atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). A top view is shown. Color indicates the distribution of deformation energy across the graphene sheet (blue corresponds to zero deformation, red corresponds to deformation with an energy of 0.008 eV per atom).
Figure 5: Stationary $l$-layer states of a nanobubble with $N_{g}=2909$ xenon atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). A top view is shown. Color indicates the distribution of pressure across the graphene sheet (blue corresponds to zero pressure, red corresponds to a pressure of $p_{n}=5$ GPa). The average pressure values are $P=1.09$, 0.90, 0.86, 0.84, and 0.85 GPa.
@@
${\bf F}_{n}=\frac{4\epsilon_{0}}{\sigma^{2}}\sum_{k=N_{c}+1}^{N}\left[2(\sigma/r_{nk})^{14}-(\sigma/r_{nk})^{8}\right]({\bf u}_{n}^{0}-{\bf u}_{k}^{0}),$
@@
Table 3: Dependence on the number of encapsulated argon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble. Values for the ground states are highlighted in italics.
@@
188 1 0 1.442 3.35 27.46 3.02 0.122
2 0.0234 1.509 5.99 25.78 2.86 0.232
260 1 0 1.357 3.35 32.23 3.11 0.084
2 0.0179 1.389 6.45 28.48 2.96 0.226
398 1 0 1.284 3.35 39.91 3.17 0.084
2 0.0086 1.240 6.53 33.90 3.05 0.193
1 0 1.194 3.36 49.24 3.26 0.068
614 2 0.0041 1.133 6.52 40.75 3.13 0.160
3 0.0145 1.151 9.07 39.78 3.06 0.228
1 0 1.131 3.35 61.63 3.30 0.054
962 2 -0.0006 1.063 6.53 49.85 3.19 0.131
3 0.0068 1.023 9.61 47.47 3.12 0.202
1 0 1.098 3.35 72.43 3.35 0.046
1335 2 -0.0033 0.982 6.53 57.02 3.24 0.115
3 0.0015 0.964 9.72 53.34 3.16 0.182
4 0.0930 0.959 12.49 51.15 3.14 0.206
1 0 1.069 3.35 84.72 3.37 0.040
1828 2 -0.0055 0.944 6.54 66.20 3.26 0.099
3 -0.0029 0.892 9.70 60.17 3.20 0.161
4 0.0005 0.941 12.41 60.66 3.17 0.205
1 0 1.038 3.35 106.65 3.41 0.031
2 -0.0080 0.872 6.54 81.66 3.31 0.080
2909 3 -0.0070 0.852 9.71 74.79 3.23 0.130
4 -0.0047 0.813 12.94 71.05 3.19 0.182
5 -0.0015 0.877 15.41 72.16 3.18 0.214
1 0 1.018 3.36 123.73 3.44 0.027
2 -0.0096 0.827 6.55 93.70 3.34 0.070
3936 3 -0.0094 0.806 9.71 85.27 3.25 0.114
4 -0.0082 0.761 12.91 79.40 3.21 0.163
5 -0.0061 0.796 15.98 78.45 3.20 0.204
The solution of problem (12) showed that each stationary state of a nanobubble always has a strictly defined number of atomic layers, $l$, in the cluster of internal atoms—see Figs. 2 and 3. For a number of encapsulated atoms $N_{g}<4000$, stationary states with $l=1,\ldots,6$ layers are possible. Each internal atom lies within one of the layers (see Fig. 3). The layers have a flat, circular shape (ranging from the shape of a regular hexagon to a circle). The layers are concentrically stacked on top of each other and form a symmetric $l$-stepped pyramid with a flat top (the plane of the layers is always parallel to the substrate plane). Atoms within each layer are in the same plane, with only edge atoms displaced downwards, smoothing the stepped shape of the pyramid. Atoms Ar, Kr, and Xe differ in size, hence the interlayer distances for them are $\Delta h=3.0$, 3.2, and 3.6 Å, respectively. For argon atoms, two- and three-layer structures at $N_{g}=2463$ were first obtained in Iakovlev2017 .
Covering the $l$-stepped pyramid from above with a graphene sheet is achieved through its local stretching. The valence bonds of the sheet stretch over the group of encapsulated atoms, while outside this zone the sheet remains in an undeformed state, lying flush against the substrate—see Fig. 4. The most significant stretching of C–C bonds occurs over the layer boundaries, where the deformation energy can reach values of 0.01 eV per carbon atom. As seen in Fig. 4, the zones of sheet stretching strictly correspond to the stepped structure of the pyramid of encapsulated atoms. For example, for $N_{g}=3936$ and $l$-layer states with $l=1,2,3$, stretching of the sheet occurs only over the edges of the layers. For states with a larger number of layers ($l=4,5$), stretching occurs more uniformly across the entire surface of the nanobubble, but the maximum values are always reached at its base (over the edge of the first layer).

### v3

@@ Abstract
Using He, Ne, Ar, Kr, and Xe atoms as a model system, it is demonstrated that graphene nanobubbles on flat substrates are multistable systems. A nanobubble can adopt multiple stable stationary states, each characterized by the number of layers $l$ within the cluster of encapsulated atoms. The layers are circular, concentrically stacked, and form an $l$-stepped pyramid with a flat top. Encapsulation of this pyramid by the graphene sheet is achieved through local stretching of the membrane: the valence bonds elongate only directly above the confined atoms. Outside this coverage zone, the sheet remains undeformed and lies flush against the substrate. The maximum number of possible layers, $l_{m}$, increases monotonically with the number of encapsulated atoms $N$, reaching $l_{m}=6$ for $N=4000$. The graphene membrane, through van der Waals interaction with the substrate, compresses the internal atomic cluster, generating pressures on the order of $P\sim 1$ GPa. Numerical simulations of thermal vibrations reveal that among all $l$-layer configurations, one ground state always exist. Upon heating, this state smoothly transitions into a layerless liquid configuration. All other stationary states transform into this ground state once a characteristic temperature $T_{l}$ is reached. For $N=4000$, the ground state corresponds to the four-layer packing ($l=4$). The coexistence of multiple stable states with distinct layer numbers at low temperatures leads to the absence of a universal shape for the nanobubbles. In this scenario, the height-to-radius ratio, $H/R$ is not constant and can vary from 0 to 0.28, depending on the number of layers.
Keywords: Van der Walls encapsulation, graphene nanobubble, molecular modelling, phase transition, aspect ratio
@@ I Introduction
Graphene nanobubbles have the potential to become indispensable tools in nanotechnology. Analysis of their shape allows for the determination of the adhesion energy between a graphene sheet and a substrate Wang2016 . Nanobubbles in multilayer graphene can be utilized in the development of cathode materials for lithium-ion batteries He2023 . Furthermore, they can be used to create two-dimensional van der Waals solids, which may find applications in quantum information technologies Langle2024 .
@@
In addition to experimental techniques, analytical models and numerical simulations have become essential tools for investigating nanobubbles. Analytical approaches are often based on elastic plate theory, which incorporates the contribution of adhesion energy between the layers of van der Waals heterostructures Yue2012 ; Wang2013a ; Dai2018 . More advanced models further integrate the equation of state of the encapsulated substance Zhilyaev2019 ; Iakovlev2019 ; Aslymov2020 ; Qu2023 and account for nonlinear effects Lyublinskaya2020 . These models have revealed that circular nanobubbles exhibit universal scaling behavior, characterized by a constant height-to-radius aspect ratio, $H/R$.
Numerical simulations serve not only to validate analytical predictions but also to provide deeper insights into the physical properties of nanobubbles. For instance, it has been shown that spatial confinement and high internal pressure can induce the formation of internal crystalline structures, which may exhibit planar geometries Iakovlev2017 . Furthermore, the equilibrium shape of a nanobubble can critically depend on the nature of the encapsulated material Ghorbanfekr2017 . Notably, simulations by S.K. Jain et al. Jain2017 demonstrated that a nanobubble containing an ideal gas invariably assumes a circular shape with a constant aspect ratio of $H/R=0.204$.
The capabilities of molecular dynamics (MD) simulations have enabled the exploration of more complex phase behaviors, such as the transition of confined argon into a liquid-crystalline state Korneva2023 . MD techniques have also been employed to model the mechanical indentation of a nanobubble by an atomic force microscope probe Faraji2022 .
From a thermal management perspective, it has been demonstrated that the presence of nanobubbles can significantly reduce the interfacial thermal conductivity of layered van der Waals structures Qu2023a . Additionally, they have been identified as a potential cause of substantial thermal contraction in these materials Qu2023b .
In this paper, all possible stationary states of graphene nanobubbles are numerically modeled. The paper is organized as follows. Section II describes the model employed. Section III presents the possible stationary states of the nanobubbles. The stability of these stationary states against thermal fluctuations is analyzed in Section IV. The main results are summarized in the Conclusion (Section V).
@@ II Model
Consider a rectangular graphene sheet in which the chains of valent bonds along the $x$-axis have a ”zigzag” structure with $N_{x}$ nodes, and along the $y$-axis—an ”armchair” structure with $N_{y}$ nodes (the sheet lies parallel to the $z=0$ plane). Consider a sheet with periodic and free boundary conditions (Fig. 1). When periodic boundary conditions are applied, the sheet consists of $N_{c}=N_{x}N_{y}/2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its periodic lengths are $L_{x}=N_{x}r_{c}\sqrt{3}/2$ and $L_{y}=3N_{y}r_{c}/4$ ($r_{c}=1.418$ Å is the C–C bond length). When free boundary conditions are applied, the sheet consists of $N_{c}=N_{x}N_{y}/2-2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its length is $L_{x}=(N_{x}-1)r_{c}\sqrt{3}/2$ and its width is $L_{y}=(3N_{y}/4-1)r_{c}$. In all calculations, we use a graphene sheet with parameters $N_{x}=416$, $N_{y}=480$ (sheet dimensions: $51.086\times 51.048$ nm2, number of carbon atoms $N_{c}=99840$). For the graphene sheet with free boundaries, we assume that hydrogen atoms are attached to the edge carbon atoms, which participate in the formation of only two C–C bonds. These edge atoms, together with the attached hydrogen atoms, form united CH atoms with a mass of $M_{e}=13m_{p}$ (the mass of the interior sheet atoms is $M_{c}=12m_{p}$, where $m_{p}=1.66\times 10^{-27}$ kg is the proton mass).
The shape of the nanobubble and its properties do not depend on the boundary conditions employed. We only note that the approach of a nanobubble to a free edge of the sheet can lead to its collapse (i.e., the release of gas atoms from under the sheet), an effect which is impossible under periodic boundary conditions.
@@
where $\cos\varphi=({\bf u}_{n}-{\bf u}_{k},{\bf u}_{l}-{\bf u}_{k})/r_{nk}r_{kl}$, and $\varphi_{0}=2\pi/3$ is the equilibrium valent angle. Parameters $\alpha=1.7889$ Å-1 and $\epsilon_{\varphi}=1.3143$ eV can be found from the small amplitude oscillations spectrum of the graphene sheet Savin2008 .
@@
where $\varepsilon_{0}$ is the interaction energy (adhesion energy) and $h_{0}$ is the equilibrium distance to the surface of the half-space. For carbon atom and silicon dioxide (SiO2) substrate, the energy is $\varepsilon_{0}=0.074$ eV and the distance is $h_{0}=5$ Å  Koenig2011 . The values used for the remaining atoms are presented in Table 1; the parameters were obtained using the Lorentz-Berthelot mixing rules.
Table 1: Values of the parameters of the interaction potential with a flat substrate (5) and masses for atoms C, He, Ne, Ar, Kr, Xe.
C He Ne Ar Kr Xe
$\varepsilon_{0}$ (eV)  0.074  0.04322  0.0797  0.1406  0.17242  0.19462
$h_{0}$ (Å) 5.0 4.577 4.662 5.0 5.116 5.397
$M_{g}$ ($m_{p}$) 12.0 4.0 20.18 39.95 83.80 131.29
@@ C He Ne Ar Kr Xe
where $N=N_{c}+N_{g}$ is the total number of atoms in the two-component molecular system, the vectors $\{{\bf u}_{n}\}_{n=N_{c}+1}^{N}$ specify the coordinates of the gas atoms, and $M_{g}$ is the atomic mass (see Table 1). The first term in the sum (6) represents the kinetic energy of the gas subsystem, the second term—the interaction energy of the gas atoms with the flat substrate. The last term,
@@
defines the interaction energy between the gas atoms, where $r_{nk}=|{\bf u}_{n}-{\bf u}_{k}|$ is the distance between atoms $n$ and $k$.
The interaction between atoms is described by the Lennard-Jones (12-6) potential:
@@
The values of the interaction potential parameters for various pairs of atoms are given in Table 2. The LJ interactions were truncated at a cutoff distance of 30 Å.
Table 2: Values of the interaction potential parameters (8) for various pairs of atoms Iakovlev2017 ; Rappe1992 ; Vogt2014 .
CHe CNe CAr CKr CXe
$\varepsilon_{1}$ (eV)  0.00161  0.00297  0.00525  0.00643  0.00724
$\sigma$ (Å) 3.0165 3.0915 3.3965 3.4965 3.7465
HeHe NeNe ArAr KrKr XeXe
$\varepsilon_{1}$ (eV) 0.00094 0.0032 0.010 0.015 0.019
$\sigma$ (Å) 2.64 2.79 3.40 3.60 4.10
@@
Figure 2: Stationary $l$-layer states of encapsulated $N_{g}=4000$ argon atoms with $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). For clarity of presentation, only one half of the graphene sheet is shown. The graphene sheet is depicted in gray. For argon atoms, their vertical displacement is indicated by color.
Figure 3: Distribution of vertical displacements $p(z)$ for the stationary states of encapsulated $N_{g}=4000$ argon atoms with number of layers $l=1$, …, 5.
@@ III Stationary States of Graphene Nanobubbles
Let $\{{\bf u}_{n}^{0}=(x_{n}^{0},y_{n}^{0},z_{n}^{0})\}_{n=1}^{N}$ denote a solution to problem (12). The corresponding stationary state is characterized by its total energy $E(\{{\bf u}_{n}^{0}\}_{n=1}^{N})$, the maximum vertical displacement of the graphene sheet above the encapsulated group of atoms (i.e., the height of the graphene bulge)
@@
Figure 4: Stationary $l$-layer states of a nanobubble with $N_{g}=4000$ krypton atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). A top view is shown. Color indicates the distribution of deformation energy across the graphene sheet.
Figure 5: Stationary $l$-layer states of a nanobubble with $N_{g}=4000$ xenon atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). A top view is shown. Color indicates the distribution of pressure across the graphene sheet. The average pressure values are $P=0.976$, 0.948, 0.977, 0.989, and 0.995 GPa.
We consider a sheet atom to participate in forming the upper surface of the nanobubble if its vertical displacement $\Delta z_{n}=z_{n}^{0}-h_{0}>h_{1}$, where the displacement threshold $h_{1}$ is 0.97, 1.17, 1.5, 1.6, and 1.8 Å  for He, Ne, Ar, Kr, and Xe atoms, respectively. The center of the nanobubble then has coordinates
@@
Table 3: Dependence on the number of encapsulated helium atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble. The number of layers in the ground state is highlighted in bold.
$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
125 1 0 1.933 2.19 13.58 0.52 0.161
250 1 0 1.641 2.12 19.86 0.55 0.107
2 0.0076 1.474 3.82 19.09 0.50 0.200
500 1 0 1.422 2.11 28.31 0.57 0.075
2 0.0049 1.227 4.33 26.09 0.51 0.166
1 0 1.252 2.10 40.54 0.59 0.052
1000 2 0.0039 1.051 4.54 36.42 0.51 0.125
3 0.0038 0.979 6.24 33.50 0.51 0.186
1 0 1.152 2.16 57.55 0.60 0.038
2 0.0026 0.930 4.45 50.40 0.52 0.088
2000 3 0.0008 0.818 6.80 44.98 0.50 0.151
1 0 1.070 2.17 92.48 0.58 0.023
2 0.0005 0.897 4.49 72.01 0.55 0.062
4000 3 -0.0014 0.758 6.80 65.84 0.51 0.103
4 -0.0025 0.694 9.01 62.92 0.49 0.143
5 -0.0031 0.652 10.44 60.11 0.49 0.174
Table 4: Dependence on the number of encapsulated neon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
125 1 0 1.736 2.52 15.81 1.69 0.159
250 1 0 1.450 2.51 22.89 1.80 0.110
2 0.0101 1.476 4.64 21.03 1.71 0.221
500 1 0 1.224 2.50 32.71 1.86 0.076
2 0.0077 1.237 4.93 28.55 1.77 0.173
1 0 1.098 2.51 46.06 1.94 0.054
1000 2 0.0034 1.007 4.92 38.17 1.83 0.129
3 0.0068 1.028 7.16 36.39 1.80 0.197
1 0 1.001 2.50 65.37 1.97 0.038
2 0.0022 0.839 4.93 52.07 1.87 0.095
2000 3 0.0034 0.872 7.40 48.46 1.83 0.153
4 0.0050 0.876 9.25 46.98 1.82 0.197
1 0 0.932 2.50 92.48 2.01 0.027
2 0.0011 0.727 4.93 72.01 1.90 0.068
4000 3 0.0015 0.760 7.38 65.84 1.85 0.112
4 0.0021 0.774 9.86 62.92 1.82 0.157
5 0.0030 0.752 11.86 60.11 1.82 0.197
@@
Here, the distance to the center is
$r_{n}=[(x_{n}^{0}-x_{c})^{2}+(y_{n}^{0}-y_{c})^{2}]^{1/2},$
and the second vertical displacement threshold $h_{2}$ is 1.62, 1.95, 2.5, 2.7, and 3.1 Å  for He, Ne, Ar, Kr, and Xe atoms, respectively.
@@
${\bf F}_{n}=\frac{24\epsilon_{0}}{\sigma^{2}}\sum_{k=N_{c}+1}^{N}\left[2(\sigma/r_{nk})^{14}-(\sigma/r_{nk})^{8}\right]({\bf u}_{n}^{0}-{\bf u}_{k}^{0}),$
@@
Table 5: Dependence on the number of encapsulated argon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
@@
125 1 0 1.599 3.48 22.31 2.80 0.156
2 0.0296 1.714 5.72 21.28 2.67 0.269
250 1 0 1.368 3.48 31.44 2.97 0.182
2 0.0179 1.480 6.57 27.61 2.87 0.238
500 1 0 1.200 3.48 44.27 3.10 0.079
2 0.0043 1.293 6.64 37.07 3.01 0.179
1 0 1.099 3.48 62.51 3.18 0.056
1000 2 -0.0052 1.148 6.64 50.02 3.12 0.133
3 0.0032 1.162 9.79 46.38 3.08 0.211
1 0 1.021 3.48 88.31 3.25 0.039
2000 2 -0.0095 1.035 6.64 68.59 3.20 0.097
3 -0.0073 1.068 9.82 62.33 3.15 0.158
4 0.0024 1.085 12.88 60.19 3.12 0.214
1 0 0.968 3.48 124.55 3.30 0.028
2 -0.0128 0.951 6.64 94.53 3.26 0.070
4000 3 -0.0134 0.997 9.82 85.73 3.20 0.115
4 -0.0116 1.019 13.02 81.56 3.17 0.160
5 -0.0080 1.025 16.09 77.39 3.16 0.208
Solving problem (12) reveals that each stationary state of a nanobubble possesses a strictly defined number of atomic layers, $l$, within the internal cluster—see Figs. 2 and 3. For systems with up to $N_{g}\leq 4000$ encapsulated atoms, stationary states with l=1,,6l=1,...,6 layers are possible. Each internal atom resides within one of these layers (Fig. 3). The layers are flat and circular in shape. They are concentrically stacked, forming a symmetric $l$-stepped pyramid with a flat top, and remain parallel to the substrate plane. While atoms within each layer lie in the same plane, the edge atoms are displaced downward, effectively smoothing the stepped profile of the pyramid. The size of the inert gas atoms varies across the series, resulting in distinct interlayer distances: $\Delta h=2.2$, 2.5, 3.0, 3.2, and 3.6 Å for He, Ne, Ar, Kr, and Xe, respectively. For argon atoms specifically, two- and three-layer structures at $N_{g}=2463$ were first reported in Iakovlev2017 .
Covering the $l$-stepped pyramid from above with a graphene sheet is achieved through its local stretching. The valence bonds of the sheet stretch over the group of encapsulated atoms, while outside this zone the sheet remains in an undeformed state, lying flush against the substrate—see Fig. 4. The most significant stretching of C–C bonds occurs over the layer boundaries, where the deformation energy can reach values of 0.01 eV per carbon atom. Here, valence bonds can elongate by 1%. This is in agreement with experimental measurements on graphene bubbles Georgiou2011 ; Zabel2011 . As seen in Fig. 4, the zones of sheet stretching strictly correspond to the stepped structure of the pyramid of encapsulated atoms. For example, for $N_{g}=4000$ and $l$-layer states with $l=1,2,3$, stretching of the sheet mainly occurs above the edges of the layers. For states with a larger number of layers ($l=4,5$), stretching occurs more uniformly across the entire surface of the nanobubble, but the maximum values are always reached at its base (over the edge of the first layer).